A bowl contains 1 green marble, 1 yellow marble, and 1 black marble. One marble is drawn from the bowl at random, then a fair die is rolled once.
Find the probability of drawing the black marble and rolling a 6.

Answer by jrfrunner(365)

P(black marble first then 6 on die) = P(Black marble)*P(6 on die) independent events
--
=(1/3)*(1/6)=1/18